OpenStudy (anonymous):

What's the equation that represents the line passing through the point (2, -4) with a slope of one half

OpenStudy (anonymous):

eq of line passing through(a,b) and having slope m is given by (y-b)=m(x-a)

OpenStudy (anonymous):

here m=1/2 and (a,b)--> (2,-4)

OpenStudy (anonymous):

thus we hv (y+4)=1/2 * (x-2)

OpenStudy (anonymous):

once you solve for the intercept you can convert it to look like y=mx+b format where m=slope and b=y-intercept